# Lost in Time: Clock and Calendar Understanding Challenges in Multimodal LLMs

### v1

@@ Appendix A Related Work
Vision-Language Benchmarks. Existing benchmarks for MLLMs cover various tasks, from college-level subject knowledge to advanced mathematical and multi-step reasoning. Massive Multi-discipline Multimodal Understanding (MMMU, Yue et al., 2024) tests deliberate reasoning across 11.5K multimodal questions drawn from college exams, quizzes, and textbooks, spanning disciplines such as Art, Business, Science, Health, Social Science, and Engineering. MathVista Lu et al. (2024) gauges mathematical reasoning within visual contexts, while GeomVerse Kazemi et al. (2024a) evaluates geometry-based problem-solving in MLLMs. ReMI Kazemi et al. (2024b) focuses on multi-image reasoning, encompassing diverse domains, including math, physics, logic, code, table/chart comprehension, and spatio-temporal tasks. Despite this breadth, none of these datasets specifically targets analogue clock and calendar interpretation. Analogue Clock and Dial Reading. Reading analogue clocks is a complex cognitive task that engages multiple mental processes Galatzer-Levy et al. (2024). It involves several key cognitive components: visuospatial skills for understanding spatial relationships between clock elements, executive functioning for planning and reasoning, working memory to maintain mental representations of time concepts, and sustained attention to process the information accurately Galatzer-Levy et al. (2024); Jones & Graff-Radford (2021). Yang et al. (2022) provide a comprehensive framework for reading analogue clocks in natural images and videos, introducing a synthetic data generation pipeline which can produce a wide variety of clock images reflecting the challenges encountered in real-world scenes. Another line of work focuses on automatic dial or gauge meter readings. The solutions proposed for dial reading rely on neural models Alexeev et al. (2020), projective transforms Bao et al. (2019), and virtual dataset generators Cai et al. (2020); Howells et al. (2021), which produce accurate results for gauges with known shape and style.
@@ Appendix B Analysis of Easy Clock Predictions
Table 2 provides model performance across different times of the day for the easy clock category. The results indicate that certain times pose greater challenges for MLLMs, with times such as 2:00, 4:00, and 5:00 being misclassified by all models. Error on this simpler task highlight gaps in accurate clock reading.

### v2

@@ 5 Conclusion
Acknowledgments
This work was supported in part by the School of Informatics at the University of Edinburgh. Rohit Saxena was supported by the EPSRC (grant no. EP/V025708/1). Aryo Pradipta Gema was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. Pasquale Minervini was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), and a donation from Accenture LLP. This work was also supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
@@ Appendix A Related Work
Vision-Language Benchmarks. Existing benchmarks for MLLMs cover various tasks, from college-level subject knowledge to advanced mathematical and multi-step reasoning. Massive Multi-discipline Multimodal Understanding (MMMU, Yue et al., 2024) tests deliberate reasoning across 11.5K multimodal questions drawn from college exams, quizzes, and textbooks, spanning disciplines such as Art, Business, Science, Health, Social Science, and Engineering. MathVista Lu et al. (2024) gauges mathematical reasoning within visual contexts, while GeomVerse Kazemi et al. (2024a) evaluates geometry-based problem-solving in MLLMs. ReMI Kazemi et al. (2024b) focuses on multi-image reasoning, encompassing diverse domains, including math, physics, logic, code, table/chart comprehension, and spatio-temporal tasks. Despite this breadth, none of these datasets specifically targets analogue clock and calendar interpretation.
Analogue Clock and Dial Reading. Reading analogue clocks is a complex cognitive task that engages multiple mental processes Galatzer-Levy et al. (2024). It involves several key cognitive components: visuospatial skills for understanding spatial relationships between clock elements, executive functioning for planning and reasoning, working memory to maintain mental representations of time concepts, and sustained attention to process the information accurately Galatzer-Levy et al. (2024); Jones & Graff-Radford (2021). Yang et al. (2022) provide a comprehensive framework for reading analogue clocks in natural images and videos, introducing a synthetic data generation pipeline which can produce a wide variety of clock images reflecting the challenges encountered in real-world scenes. Recently, Deitke et al. (2024) introduced a large-scale synthetic dataset of analogue clocks comprising images rendered from different watch models. However, the dataset’s generation pipeline is not publicly available, making it difficult to reproduce. Ghosal et al. (2024) propose a puzzle-based task that includes clock puzzles to identify when an event occurred or will occur, given a starting time and an elapsed or future duration. Another line of work focuses on automatic dial or gauge meter readings. The solutions proposed for dial reading rely on neural models Alexeev et al. (2020), projective transforms Bao et al. (2019), and virtual dataset generators Cai et al. (2020); Howells et al. (2021), which produce accurate results for gauges with known shape and style.
@@ Appendix B Analysis of Easy Clock Predictions
Table 2 provides model performance across different times of the day for the easy clock category. The results indicate that certain times pose greater challenges for MLLMs, with times such as 2:00, 4:00, and 5:00 being misclassified by all models. Errors on this simpler task highlight gaps in accurate clock reading.